# Deep Generative Models for Detector Signature Simulation: An Analytical Taxonomy

### v1

@@ Abstract
In modern collider experiments, the quest to explore fundamental interactions between elementary particles has reached unparalleled levels of precision. Signatures from particle physics detectors are low-level objects encoding the physics of collisions. The complete simulation of them in a detector is a memory and storage-intensive task. To address this computational bottleneck in particle physics, “Fast Simulation” has been introduced and refined over the years. The field has seen a surge in interest in surrogate modeling the detector simulation, fueled by the advancements in deep generative models. These models aim to generate responses that are statistically identical to the observed data. In this paper, we conduct a comprehensive and exhaustive taxonomic review of the existing literature on the simulation of detector signatures from both methodological and application-wise perspectives. Initially, we formulate the problem of detector signature simulation and discuss its different variations that can be unified. Next, we classify the state-of-the-art methods into four distinct categories based on their underlying model architectures, summarizing their respective generation strategies. We then identify and discuss three key application areas. Finally, we shed light on the challenges and opportunities that lie ahead in detector signature simulation, setting the stage for future research and development.
@@ 1 Introduction
The core concept of generative models is derived from the training of a density estimator that produces samples approximating the distribution of the training data. The initial generation of neural network-based generative models, also known as energy-based models [1, 2], attempted to accomplish this by establishing an energy function proportional to the likelihood for data points. However, these models faced challenges in scaling up to high-dimensional, complex data, like natural images. Thus, they necessitated the use of Markov Chain Monte Carlo (MCMC) sampling [3], a method required during both training and inference stages. The goal of energy-based models is to adjust the model parameters so that the energy function assigns lower energy (i.e., higher likelihood) to real data points and higher energy to improbable ones. However, the challenge is that these energy landscapes can be very complex, especially for high-dimensional data. Calculating likelihoods or gradients directly becomes computationally infeasible or extremely difficult. That’s where MCMC comes in. This method, characterized by its iterative nature, often resulted in a very slow and inefficient process. The resurgence of interest in generative models in the past few years can be attributed to the availability of larger datasets and significant advancements in Deep Generative Models (DGMs) [4] for natural images. These DGMs have pushed the boundaries in terms of visual quality, sample diversity, and speed of sampling.
@@
In Particle Physics, the application of DGMs was first studied in the “Fast and Efficient Simulation” campaign [5] that sparked the search for faster and more storage-efficient surrogate models and simulation methods of collider physics experiments [6, 7]. Surrogate models for fast detector simulation are simplified, computationally efficient approximations that emulate the behavior of more complex, detailed simulations of signatures from particle detectors. These models are constructed using DGMs, trained on a dataset either generated by the Geant4 simulation [8, 9, 10] or the real signatures from particle detectors. Once trained, the model can generate samples that are statistically similar to their training data and even can generalize to the data beyond the training regime [11, 12, 13, 14] depending on the objective. There are some key requirements that the model should meet to be effective and efficient:
@@
    Low time-complexity: The surrogate model must be computationally efficient in order to facilitate fast simulations. Computational speed is crucial when performing large-scale simulations or when needing to iterate the model many times for optimization or fine-tuning. Thus, the model should take advantage of parallel processing capabilities and have to be optimized for the hardware it is expected to run on, whether it’s a CPU or GPU.
@@
    Low space-complexity: It should come with a minimal storage cost. Hence, the underlying compression technique has to reduce the storage footprint without sacrificing too much in terms of the downstream physics analysis.
@@
The subsequent sections of this paper are organized in the following manner. We begin by defining the challenge of simulating signatures from particle detectors with deep generative models as surrogates. Subsequently, we present a taxonomic breakdown of current methodologies, categorizing them into four classes. We introduce a universal framework, delve into prevalent generation strategies and tasks, and then provide an exhaustive review of research for each category. In our concluding remarks, we address the challenges and potential advancements in deep generative modeling for detector response generation. Through this survey, we aim to offer a unified and consolidated view of the present landscape of deep generative models for detector signature simulation. We hope to bridge the existing gap in this sparse and occasionally overlooked research, ensuring an inter-experimental perspective that remains objective and free from bias.
@@ 2 Problem Definition
The detector signatures can be defined by a triplet $\mathcal{D}_{e}=(\mathbf{L}_{e},\mathbf{C}_{e},\mathbf{H}_{e})$ for each event $e$, where $\mathbf{L}\in\mathbb{Z}$ is the detector component such as layers or sensors indicators, $\mathbf{C}\in\mathbb{R}^{n}$ is the global attribute of the corresponding event such as incident energy or the beam parameter, $\mathbf{H}\in\mathbb{R}^{d}$ is the hit points per sensor/layer like tracks, clusters or jets which can be represented for instance by a grid-like an image and sequence, or a multi-set like point clouds and graphs. Given a set of $M$ observed detector responses $\mathbf{D}=\{\mathcal{D}^{i}_{e}\}_{i=1}^{M}$, DGMs learn the distribution of these signatures $p(\mathbf{D})$, from which new responses can be sampled $\mathbf{H}_{\text{new}}\sim p(\mathbf{H})$.
This paper presents both the latent variable and non-latent variable approaches as the mainstream detector response generation models. Latent variable approaches follow an encoder–sampler–decoder pipeline. It firstly maps the data into a hidden space through an encoding function, manipulates the hidden variables to reflect the desired properties of the detector response to be generated, and then generates new samples based on latent codes through a decoding function. Unlike the latent variable models111 Here, the term latent refers to a compressed space. In the literature, latent spaces sometimes also refer to uncompressed spaces, such as for normalizing flows or diffusion models., non-latent variable approaches directly map the input data to the desired detector responses stochastically without the intermediary step of encoding into a compressed hidden space.
@@ 3 Algorithm Taxonomy
The Encoder. The encoding function $f_{\theta}(z|\mathcal{D})$ maps the detector response triplets to a dense, continuous (or quantized) topological space. To ensure the learned latent space is meaningful for generation, depending on the data type and the inductive bias, one incorporates various morphisms (e.g., Convolutional Neural Networks [15], Graph Neural Networks [16], DeepSets [17], etc) as the encoder. In this way, the encoder function $f_{\theta}$ generates the parameters for a stochastic distribution that adheres to a prior distribution denoted as $p(z)$.
The Sampler. Following the encoding process, the model that generates the detector response draws samples from the latent distribution $z\sim p(z)$. There are generally two prevalent strategies for this sampling process: random sampling and controlled sampling [18]. Random sampling refers to randomly selecting latent vectors from the learned or prior distribution. On the other hand, guided or controlled sampling is designed to sample the stratified latent vectors with the specific goal of generating samples that exhibit certain preferred characteristics. These characteristics could be certain kinematic profiles, beam parameters, physics processes, number of hits, luminosities, incident angles, or incident energies. In most tasks, controlled sampling is model-dependent and necessitates an additional optimization component that goes beyond the scope of random generation.
The Decoder. Upon obtaining the latent features drawn from the learned distribution, the decoding mechanism is responsible for reconstituting them into a data manifold. Due to the multi-objective and fine-grained learning nature of detector simulation, the decoding phase is inherently more complex than the encoding stage. Typically, the decoders could be grouped into three categories: sequential generation, one-shot generation, and zero-shot generation. Sequential generation refers to generating the detector component information in a set of consecutive and autoregressive steps, usually done sensor-by-sensor, layer-by-layer, or hit-by-hit. One-shot generation, instead, refers to generating the whole detector signatures in one single step. Zero-shot generation is when the model generates an unseen set of detector information where the control parameter goes beyond the training data. Thus, the model has to generalize well to the Out-Of-Distribution (OOD) domain. While zero-shot generation may employ one-shot or sequential methods for the actual data construction, its main challenge lies in ensuring that these generated responses are both accurate and plausible when dealing with conditions that differ from the training data.
@@ 3.1 Deep Generative Models
Figure 1: Generic architectures of deep Generative models for simulating detector signatures
Variational Auto Encoders (VAE). VAE estimates the distributions of data $p(\mathcal{D})$ with the amortized variational posterior [19, 20], by maximizing the Evidence Lower BOund (ELBO), i.e. the lower bound of the log-likelihood function, as follows:
@@
where the former term is known as the negative reconstruction loss between the input and the reconstructed data, while the latter is the disentanglement enhancement regularizer that drives the amortized variational posterior $f_{\theta}(z|\mathcal{D})$ to the prior (marginal) distribution $p_{\theta}(z)$. For simulation-based inference where the tractability of the likelihood is vital, one can use various sampling methods to estimate the likelihood such as Importance Sampling [20, 21], Annealed Importance Sampling [22, 23], and Nested Variational Inference [24] with a sufficiently large sample space to estimate the log-likelihood. Generally, VAEs require fewer computational resources than all the DGMs, converge faster, and one can have more control over the conditional parameters.
However, VAEs face a few challenges. A potential issue is the “posterior collapse” [25], which occurs when a powerful decoder treats the latent variable $z$ as mere noise, leading to the regularization term being minimized for priors like the standard Gaussian. Another problem, known as the “hole problem” [26], arises from a mismatch between the aggregated posterior and the prior. If there are regions where the prior assigns high probability, but the aggregated posterior assigns low, sampling from these regions can result in low-quality output.
Normalizing Flows. Normalizing flow (NF) estimates the density of data $p(\mathcal{D})$ directly with an invertible and deterministic bijection between the latent variables and the data manifold via the change of variables formula. A typical instance of flow-based models takes the following form:
@@
where $p(\bm{z})$ is the density of the latent variables $\bm{z}$. The term $f^{-1}(\mathcal{D})$ is the inverse of the transformation $f$ that maps the latent variables to the data. The determinant of the Jacobian matrix $\left|\det\left(\frac{\partial f^{-1}(\mathcal{D})}{\partial\mathcal{D}}\right)\right|$ accounts for the change in volume induced by this transformation, allowing for efficient and expressive density estimation. Flow-based models are proper candidates for lossless compression [27, 28], and Approximate Bayesian Computation (ABC) [29] in simulation-based inference since they allow the calculation of the exact likelihood. However, the likelihood does not always correlate directly with sample quality [30]. Conditional flow-based models could also be used to form a flexible family of variational posteriors [31, 32] in VAE-based models with which the lower bound to the log-likelihood function could be tighter.
The main problem with NF-based models is the computational overhead they create with high-dimensional data. NF models can struggle with very high-dimensional data spaces as the complexity of the necessary transformations grows exponentially. This makes them less suitable for simulating complex and high granular detector signatures like HGCAL [33] for HL-LHC [34] and PXD [35] at Belle II [36].
Generative Adversarial Networks (GAN). A GAN [37, 38] is an implicit density estimator [39], which learns to sample real data points in a zero-sum game. GANs consist of two main components, namely, a generator $f_{G}$ for estimating the density implicitly and generating realistic data and a discriminator $f_{D}$ for distinguishing between the generated and real detector responses. This adversarial mechanism encourages the generator to produce synthetic data that the discriminator cannot distinguish from real data, improving the model’s ability to capture the data distribution. Although GANs do not learn the density directly, they do it implicitly, where one can indeed incorporate it for simulation-based inference [40], for example, in detector unfolding [41, 42]. Formally, its training objective is a min-max game as follows:
@@
In general, training GANs is a highly brittle and painful task. It requires a significant amount of patience and hyperparameter tuning for domain-specific tasks, but when it generalizes well, it can produce remarkably high-resolution samples [43, 44] with high fidelity. As with almost all DGMs, GANs also suffer from mode collapse and vanishing gradients. Nevertheless, GANs provide good control over the conditioning and enable representation learning [45] for any possible downstream task. However, a specific issue with GANs is their evasive Nash equilibrium, a state where the discriminator and the generator cannot unilaterally improve their position by deviating from their current strategy. The models may continue to oscillate without settling into a stable state, or they may converge to a non-optimal solution. This makes the model selection non-trivial, since GANs have no straightforward way to select the best training epoch.
@@
Autoregressive Models Autoregressive Models (ARMs) [46] are based on the chain rule of probability and decompose a joint distribution over $N$ random variables. Specifically, this model factorizes the generation process as a sequential step, which determines the next step action given an initial detector layer, hit, jet constituent, or shower. The general formulation of ARM models is as follows:
$p(\mathcal{D}^{e})=\prod_{i=1}^{N}p(\mathcal{D}^{e}_{i}\mid\mathcal{D}^{e}_{1},\mathcal{D}^{e}_{2},\cdots,\mathcal{D}^{e}_{i-1}),$ (4)
@@
where it is possible to directly maximize the likelihood of the data by training a recurrent neural network to model $p(\mathcal{D}^{e}_{i}\mid\mathcal{D}^{e}_{1:i-1})$ by minimizing the negative log-likelihood,
$-\ln p(\mathcal{D}^{e})=-\sum_{i}\ln p(\mathcal{D}^{e}_{i}\mid\mathcal{D}^{e}_{1},\mathcal{D}^{e}_{2},\cdots,\mathcal{D}^{e}_{i-1}).$ (5)
@@
Since ARMs work like sequential generation, applying AR models requires a pre-specified ordering defined by hit orders $\mathbf{H}_{e}$ or sensor/layer $\mathbf{L}_{e}$ orders in a detector signature triplet. While autoregressive models are extremely powerful density estimators, sampling is inherently a sequential process and can be exceedingly slow on high-dimensional data. Additionally, data must be decomposed into a fixed ordering (e.g., temporal); while the choice of ordering can be clear for detector layers in regular detectors, it is not obvious for irregular and complex topologies, which can affect performance.
@@
Diffusion Models Diffusion models or deep score-based [47, 48] generative models are a class of generative processes inspired by non-equilibrium thermodynamics [49] and PDEs [50]. Score matching is based on the idea of minimizing the difference between the derivatives of the data and the model’s log density functions. Diffusion models contain two processes, the forward and the reverse diffusion process. The forward diffusion process constantly adds noise to the data sample using noise schedule $\beta_{t}\in(0,1)$ controlling the step size, while the reverse diffusion process recreates the true data sample from a Gaussian noise input. In the language of hierarchical VAEs [51, 52], the bottom-up path (i.e., the variational posteriors) can be a diffusion process, and the top-down path is parameterized by a reversed diffusion. Ergo, they provide only an approximation of the likelihood. Formally, the forward diffusion process from step $(t-1)$ to $t$ is defined as a discrete Markov chain

### v2

# urldate  [^2] The author changed his first name from Hosein to Baran

@@ Abstract
In modern collider experiments, the quest to explore fundamental interactions between elementary particles has reached unparalleled levels of precision. Signatures from particle physics detectors are low-level objects (such as energy depositions or tracks) encoding the physics of collisions (the final state particles of hard scattering interactions). The complete simulation of them in a detector is a computational and storage-intensive task. To address this computational bottleneck in particle physics, alternative approaches have been developed, introducing additional assumptions and trade off accuracy for speed. The field has seen a surge in interest in surrogate modeling the detector simulation, fueled by the advancements in deep generative models. These models aim to generate responses that are statistically identical to the observed data. In this paper, we conduct a comprehensive and exhaustive taxonomic review of the existing literature on the simulation of detector signatures from both methodological and application-wise perspectives. Initially, we formulate the problem of detector signature simulation and discuss its different variations that can be unified. Next, we classify the state-of-the-art methods into five distinct categories based on their underlying model architectures, summarizing their respective generation strategies. Finally, we shed light on the challenges and opportunities that lie ahead in detector signature simulation, setting the stage for future research and development.
@@ 1 Introduction
Simulation is at the core of research in experimental particle physics. It allows for a detailed comparison of experimental data and fundamental theory, and therefore for a proper interpretation of measurements. Following the hierarchy of the involved energy scales, the simulation chain is split into the simulation of the hard (high-energy) scattering interaction; the decay of the heavy, unstable particles and the evolution of a parton shower; the hadronization of all colored particles to color-neutral ones; and finally, the interaction of these stable particles with the detectors [butter_machine_2023, campbell_event_2024]. In this review, we will focus on the last step, how the signatures of particles with given initial conditions (like sensor positions, particle type or energy) in the detector look like, as this is usually the computationally most-expensive step. The most detailed and accurate simulation toolboxes at hand for this task are Geant4 [agostinelli_geant4simulation_2003, allison_geant4_2006, noauthor_recent_nodate] and FLUKA [ferrari_fluka_2005, noauthor_fluka_nodate]. These track each particle and all the secondary particles it produces via fundamental interactions through all different types of matter in the entire available geometry. These simulations result in a wide variety of representations for different types of detector signatures. Among them, the showers of particles in calorimeters; the collimated, narrow cones of hadrons (called jets) produced during the hadronization of colored partons; and hit patterns that form tracks of charged particles — all yielding datasets with different granularity, complexity and characteristics.
However, the complexity and accuracy of these simulations come at the price of being computationally expensive, easily becoming the bottleneck in the simulation chain. This led to the development of several faster alternatives, trading accuracy for speed. For example, each experimental collaboration developed their own so-called “Fast Simulation”, a dedicated detector simulation making several simplifying assumptions on both, geometry and physics, leading to a speed-up of $\mathcal{O}(10)$$\mathcal{O}(100)$ [abdullin_fast_2011, atlas_collaboration_atlfast3_2022]. If one is only interested in few, high-level observables like isolated leptons or missing transverse energy, one can resort to Delphes [de_favereau_delphes_2014, selvaggi_delphes_2014, mertens_new_2015]. It is extensively used in phenomenological studies, especially also since experiment-specific Fast Simulations are restricted to members of the respective collaborations. By including even the reconstruction steps, one achieves a very fast end-to-end simulator called FlashSim [vaselli_flashsim_2023]. The common denominator of all of these approaches is that the underlying physics is stochastic: Given the same initial conditions, the resulting recorded patterns in the detector are random samples drawn from a (complicated) distribution.
@@
This is where Deep Generative Models (DGMs) will be useful. The core concept of generative models is derived from the training of a density estimator that produces samples approximating the distribution of the training data. The essence of generative models is rooted in the domain of density estimation [parzen_estimation_1962, b_w_silverman_density_1998], where various methods [fukunaga_optimization_1973, magdon-ismail_neural_1998, cranmer_kernel_2001], both parametric and non-parametric approaches, were employed to estimate the underlying distribution of data. The initial wave of neural network-based generative models, also known as energy-based models [hinton_optimal_1983, hinton_training_2002], attempted to accomplish this by establishing an energy function proportional to the likelihood for data points. However, these models faced challenges in scaling up to high-dimensional, complex data, like natural images. Thus, they necessitated the use of Markov Chain Monte Carlo (MCMC) sampling [hinton_fast_2006], a method required during both training and inference stages. The goal of energy-based models is to adjust the model parameters so that the energy function assigns lower energy (i.e., higher likelihood) to real data points and higher energy to improbable ones. However, the challenge is that these energy landscapes can be very complex, especially for high-dimensional data. Calculating likelihoods or gradients directly becomes computationally infeasible or extremely difficult. That is where MCMC comes in. This method, characterized by its iterative nature, often resulted in a very slow and inefficient process. The availability of larger datasets and significant advancements in DGMs [tomczak_deep_2022] for natural images led to a resurgence of interest in generative models in the past few years. These DGMs have pushed the boundaries in terms of visual quality, sample diversity, and speed of sampling.
@@
In Particle Physics, the application of DGMs was first studied in the “Fast and Efficient Simulation” campaign [de_oliveira_learning_2017] that sparked the search for faster and more storage-efficient surrogate models and simulation methods of collider physics experiments [adelmann_new_2022, butter_machine_2023]. Surrogate models for fast detector simulation are simplified, computationally efficient approximations that emulate the behavior of more complex, detailed simulations of signatures from particle detectors. These models are constructed using DGMs, trained on a dataset either generated by the Geant4 simulation [agostinelli_geant4simulation_2003, allison_geant4_2006, noauthor_recent_nodate] or the real signatures from particle detectors. Once trained, the model can generate samples that are statistically similar to their training data and even can generalize to the data beyond the training regime [butter_ganplifying_2021, bieringer_calomplification_2022, matchev_uncertainties_2022, hashemi_deep_2024] depending on the objective. There are some key requirements that the model should meet to be effective and efficient:
@@
    Low time-complexity: The surrogate model must be computationally efficient in order to facilitate Fast Simulations. Computational speed is crucial when performing large-scale simulations or when needing to iterate the model many times for optimization or fine-tuning. Thus, the model should take advantage of parallel processing capabilities and have to be optimized for the hardware it is expected to run on, whether it is a CPU or GPU.
@@
    Low space-complexity: It should come with a minimal storage cost. Hence, the underlying compression technique has to reduce the storage footprint without sacrificing too much of the precision and accuracy of the downstream physics analysis.
@@
The subsequent sections of this paper are organized in the following manner. We begin by defining the challenge of simulating signatures from particle detectors with deep generative models as surrogates. Subsequently, we present a taxonomic breakdown of current methodologies, categorizing them into five classes. We introduce a universal framework, delve into prevalent generation strategies and tasks, and then provide an exhaustive review of research for each category. In our concluding remarks, we address the challenges and potential advancements in deep generative modeling for detector response generation. Through this survey, we aim to offer a unified and consolidated view of the present landscape of deep generative models for detector signature simulation. We hope to bridge the existing gap in this sparse and occasionally overlooked research, ensuring an inter-experimental perspective that remains objective and free from bias.
@@ 2 Problem Definition
The detector signatures can be defined by a triplet $\mathcal{D}_{e}=(\mathbf{L}_{e},\mathbf{C}_{e},\mathbf{H}_{e})$ for each event $e$, where $\mathbf{L}\in\mathbb{Z}$ is the detector component such as layers or sensors indicators, $\mathbf{C}\in\mathbb{R}^{n}$ is the global attribute of the corresponding event such as incident energy or the beam parameter, $\mathbf{H}\in\mathbb{R}^{d}$ is the hit points per sensor/layer like particle tracks, energy clusters or jets, which can be represented for instance by a grid (like an image), by a sequence, or by a multi-set (like point clouds and graphs). Given a set of $M$ observed detector responses $\mathbf{D}=\{\mathcal{D}^{i}_{e}\}_{i=1}^{M}$, DGMs learn the distribution of these signatures $p(\mathbf{D})$, from which new responses can be sampled $\mathbf{H}_{\text{new}}\sim p(\mathbf{H})$.
In the context of detector simulation for physics analyses, $\mathbf{C}$ would be defined implicitly by the underlying hard scattering process and subsequent shower and decays, yielding a distribution of incident parameters for each particle (per event) to be simulated. The direction of motion of these particles also implicitly defines which component of the detector $\mathbf{L}$ will be relevant for the simulation. In the context of working with DGMs, these parameters are usually sampled from a convenient, externally-given distribution (for example log-uniform incident energies) or even fixed completely (for example by considering only particles traveling perpendicular to the detector surface).
This paper presents both latent variable and non-latent variable approaches as the mainstream detector response generation models. Latent variable approaches follow an encoder–sampler–decoder pipeline. It firstly maps the data into a hidden (latent) space through an encoding function, manipulates the hidden variables to reflect the desired properties of the detector response to be generated, and then generates new samples based on latent codes through a decoding function. Unlike the latent variable models333 Here, the term latent refers to a compressed space. In the literature, latent spaces sometimes also refer to uncompressed spaces, such as for normalizing flows or diffusion models., non-latent variable approaches directly map the input data to the desired detector responses stochastically without the intermediary step of encoding into a compressed hidden space.
@@ 3 Algorithm Taxonomy
The Encoder. The encoding function $f_{\theta}(z|\mathcal{D})$ maps the detector response triplets to a dense, continuous (or quantized) topological space. To ensure the learned latent space is meaningful for generation, depending on the data type and the inductive bias, one incorporates various morphisms (e.g., Convolutional Neural Networks [fukushima_neocognitron_1980], Graph Neural Networks [velickovic_everything_2023], DeepSets [zaheer_deep_2018], etc.) as the encoder. In this way, the encoder function $f_{\theta}$ generates the parameters for a stochastic distribution that adheres to a prior distribution denoted as $p(z)$.
The Sampler. Following the encoding process, the model that generates the detector response draws samples from the latent distribution $z\sim p(z)$. There are generally two prevalent strategies for this sampling process: random sampling and controlled sampling [goodman_controlled_1950]. Random sampling refers to randomly selecting latent vectors from the learned or prior distribution. On the other hand, guided or controlled sampling is designed to sample the stratified latent vectors with the specific goal of generating samples that exhibit certain preferred characteristics. These characteristics could be certain kinematic profiles, beam parameters, physics processes, number of hits, luminosities, incident angles, or incident energies. In most tasks, controlled sampling is model-dependent and necessitates an additional optimization component that goes beyond the scope of random generation.
The Decoder. Upon obtaining the latent features drawn from the learned distribution, the decoding mechanism is responsible for reconstituting them into a data manifold. Due to the multi-objective and fine-grained learning nature of detector simulation, the decoding phase is inherently more complex than the encoding stage. In other words, it is of paramount importance that the underlying symmetries/inductive biases of the detector signature is satisfied at the decoding stage. Typically, the decoders could be grouped into three categories: sequential generation, one-shot generation, and zero-shot generation. Sequential generation refers to generating the detector component information in a set of consecutive and autoregressive steps, usually done sensor-by-sensor, layer-by-layer, or hit-by-hit. One-shot generation, instead, refers to generating the whole detector signatures in one single step. Zero-shot generation is when the model generates an unseen set of detector information where the control parameter goes beyond the training data. Thus, the model has to generalize well to the Out-Of-Distribution (OOD) domain. While zero-shot generation may employ one-shot or sequential methods for the actual data construction, its main challenge lies in ensuring that these generated responses are both accurate and plausible when dealing with conditions that differ from the training data.
@@ 3.1 Deep Generative Models
Figure 1: Generic architectures of deep Generative models for simulating detector signatures.
Variational Auto Encoders (VAE). VAEs learn a lower-dimensional representation of the data in the latent space by forcing the information through a bottleneck. A VAE estimates the distributions of data $p(\mathcal{D})$ with the amortized variational posterior [kingma_auto-encoding_2022, rezende_stochastic_2014], by maximizing the Evidence Lower BOund (ELBO), i.e. the lower bound of the log-likelihood function, as follows:
@@
where the former term is known as the negative reconstruction loss between the input and the reconstructed data, while the latter is the disentanglement enhancement regularizer that drives the amortized variational posterior $f_{\theta}(z|\mathcal{D})$ to the prior (marginal) distribution $p_{\theta}(z)$. For simulation-based inference where the tractability of the likelihood is vital, one can use various sampling methods to estimate the likelihood such as Importance Sampling [rezende_stochastic_2014, burda_importance_2016], Annealed Importance Sampling [neal_annealed_2001, zhang_differentiable_2021], and Nested Variational Inference [zimmermann_nested_2021] with a sufficiently large sample space to estimate the log-likelihood. Generally, VAEs require fewer computational resources than all the DGMs, converge faster, and one can have more control over the conditional parameters. They could be adopted for controllable sampling by either modifying the loss function to enforce latent variables to be correlated with properties of interest or to feed the conditional information to different parts of the model. VAEs can also be incorporated to compress data in hybrid DGMs. However, VAEs face a few challenges [kingma_introduction_2019, tomczak_deep_2022]. A potential issue is the “posterior collapse” [bowman_generating_2016], which occurs when a powerful decoder treats the latent variable $z$ as mere noise, leading to the regularization term being minimized for priors like the standard Gaussian. Another problem, known as the “hole problem” [rezende_taming_2018], arises from a mismatch between the aggregated posterior and the prior. If there are regions where the prior assigns high probability, but the aggregated posterior assigns low, sampling from these regions can result in low-quality output. Moreover, minimizing the KL divergence between the approximated and true posterior distributions, can lead to a variance mismatch when these distributions are not perfectly aligned. This issue is known as “blurriness problem” [kingma_introduction_2019] and is addressed by using more flexible inference and generative models to better capture the data distribution and improve sample fidelity.
Normalizing Flows. A normalizing flow (NF) estimates the density of data $p(\mathcal{D})$ directly with an invertible and deterministic bijection between the latent variables and the data manifold via the change of variables formula. A typical instance of flow-based models takes the following form:
@@
where $p(\bm{z})$ is the density of the latent variables $\bm{z}$. The term $f^{-1}(\mathcal{D})$ is the inverse of the transformation $f$ that maps the latent variables to the data. The determinant of the Jacobian matrix $\left|\det\left(\frac{\partial f^{-1}(\mathcal{D})}{\partial\mathcal{D}}\right)\right|$ accounts for the change in volume induced by this transformation, allowing for efficient and expressive density estimation. Flow-based models are proper candidates for lossless compression [hoogeboom_integer_2019, berg_idf_2021], and Approximate Bayesian Computation (ABC) [papamakarios_sequential_2019] in simulation-based inference since they allow the calculation of the exact likelihood. However, the likelihood does not always correlate directly with sample quality [theis_note_2016]. Conditional flow-based models could also be used to form a flexible family of variational posteriors [rezende_variational_2016, berg_sylvester_2019] in VAE-based models with which the lower bound to the log-likelihood function could be tighter.
The main problem with NF-based models is the computational overhead they create with high-dimensional data. NF models can struggle with very high-dimensional data spaces as the complexity of the necessary transformations grows exponentially. This makes them less suitable for simulating complex and high granular detector signatures like the planned CMS High Granularity Calorimeter (HGCAL) [magnan_hgcal_2017] for HL-LHC [bruning_chapter_2020] and the Pixel Vertex Detector (PXD) [giakoustidis_status_2023] at Belle II [abe_belle_2010].
Generative Adversarial Networks (GAN). A GAN [schmidhuber_making_1990, goodfellow_generative_2014] is an implicit density estimator [noauthor_statistics_nodate], which learns to sample data points in a zero-sum game. GANs consist of two main components, namely, a generator $f_{G}$ for estimating the density implicitly and generating realistic data and a discriminator $f_{D}$ for distinguishing between the generated and real detector responses. This adversarial mechanism encourages the generator to produce synthetic data that the discriminator cannot distinguish from real data, improving the model’s ability to capture the data distribution. Although GANs do not learn the density directly, they do it implicitly, where one can indeed incorporate it for simulation-based inference [ramesh_gatsbi_2022], for example, in detector unfolding [datta_unfolding_2018, bellagente_how_2020]. Formally, its training objective is a min–max game as follows:
@@
GAN-based models, by design, allow easy implementation of controllable sampling due to introducing a sample discriminator given the desired properties, though taming its convergence is very difficult. In general, training GANs is a highly brittle and painful task. It requires a significant amount of patience and hyperparameter tuning for domain-specific tasks, but when it generalizes well, it can produce remarkably high-resolution samples [sauer_stylegan-xl_2022, noauthor_ffhq_nodate] with high fidelity. The adversarial training makes GANs suffer from mode collapse and vanishing gradients. That is why during the past few years many extensions and auxiliary methods have been introduced to mitigate these issues [salimans_improved_2016, kodali_convergence_2017, arjovsky_wasserstein_2017, bellemare_cramer_2017, miyato_spectral_2018, brock_large_2019, karras_style-based_2019, cont_tail-gan_2023]. Nevertheless, GANs provide good control over the conditioning and enable representation learning [kang_contragan_2021, hashemi_ultra-high-resolution_2023]. Representation learning refers to deriving meaningful and often stronger latent features from raw data that can be used effectively in various downstream tasks such as classification or prediction. However, a specific issue with GANs is their evasive Nash equilibrium, a state where the discriminator and the generator cannot unilaterally improve their position by deviating from their current strategy. The models may continue to oscillate without settling into a stable state, or they may converge to a non-optimal solution. This makes the model selection non-trivial, since GANs have no straightforward way to select the best training epoch.
@@
Autoregressive Models Autoregressive Models (ARMs) [bengio_neural_2000, oord_conditional_2016, chen_pixelsnail_2018, you_graphrnn_nodate] are based on the chain rule of probability and decompose a joint distribution over $N$ random variables. They are designed to generate data in a sequential manner. Specifically, these model factorize the generation process as a sequential step, which determines the next step action given an initial detector layer, hit, jet constituent, or shower. The general formulation of ARM models is as follows:
$p(\mathcal{D}_{e})=\prod_{i=1}^{N}p(\mathcal{D}_{e}^{i}\mid\mathcal{D}_{e}^{1},\mathcal{D}_{e}^{2},\cdots,\mathcal{D}_{e}^{i-1}),$ (4)
@@
where it is possible to directly maximize the likelihood of the data by training a recurrent neural network to model $p(\mathcal{D}_{e}^{i}\mid\mathcal{D}_{e}^{1:i-1})$ by minimizing the negative log-likelihood,
$-\ln p(\mathcal{D}_{e})=-\sum_{i}\ln p(\mathcal{D}_{e}^{i}\mid\mathcal{D}_{e}^{1},\mathcal{D}_{e}^{2},\cdots,\mathcal{D}_{e}^{i-1}).$ (5)
@@
Although these models might utilize internal hidden states during the generation process, it is worth noting that these states are not considered latent variables in the traditional probabilistic sense. Since ARMs work like sequential generation, applying these models requires a pre-specified ordering defined by hit orders $\mathbf{H}_{e}$ or sensor/layer $\mathbf{L}_{e}$ orders in a detector signature triplet. Thus, they offer the advantage of generating highly correlated data, thanks to their sequential conditioning on previously generated elements. This enables intricate probabilistic relationships to be captured, offering the potential for more realistic simulations. While autoregressive models are extremely powerful density estimators, sampling is inherently a sequential process and can be exceedingly slow on high-dimensional data. Additionally, data must be decomposed into a fixed ordering (e.g., temporal); while the choice of ordering can be clear for detector layers in regular detectors, it is not obvious for irregular and complex topologies, which in turn can affect performance. Hashemi et al. [hashemi_deep_2024] tried to solve this problem with proper a positional embedding and inductive bias injection.
@@
Diffusion Models Diffusion models or deep score-based generative models  [song_generative_2020, ho_denoising_2020, song_how_2021, kingma_variational_2023, ramesh_hierarchical_2022] are a class of generative processes inspired by non-equilibrium thermodynamics [sohl-dickstein_deep_2015] and PDEs [tzen_neural_2019]. Score matching is based on the idea of minimizing the difference between the derivatives of the data and the model’s log density functions. Diffusion models contain two processes, the forward and the reverse diffusion process. The forward diffusion process constantly adds noise to the data sample using a noise schedule $\beta_{t}\in(0,1)$ controlling the step size, while the reverse diffusion process recreates the true data sample from a Gaussian noise input. In the language of hierarchical VAEs [kingma_improved_2016, sonderby_ladder_2016], the bottom-up path (i.e., the variational posteriors) can be a diffusion process, and the top-down path is parameterized by a reversed diffusion. Ergo, they provide only an approximation of the likelihood. Formally, the forward diffusion process from step $(t-1)$ to $t$ is defined as a discrete Markov chain